OpenStudy (anonymous):

What is the solution for: y=x+3 y=2x

mathslover (mathslover):

It is given to you in the second equation that \(\sf{y = 2x}\) . So, just put y = 2x in the first equation : y = x + 3 As y = 2x(from eq. 2) So, eq.1 becomes : 2x = x + 3

mathslover (mathslover):

Now, solve for "x" and then for "y" . Can you do that?

OpenStudy (anonymous):

2x=x+3 2x-x=3 x=3 Answer: 3